7.

7) Transpose the following determinant: Is det(A) = det(A)? Answers: A = cl, yes det(A) = det(AT) = 24.
Answer #1

solution 7 A= 13 -61 116 det (A) = ad-bc +(3)(6) -(-6)(4) - 18-(-6) det (A) = 24 det(AT) = (3) (6)-(1)(-6) = 18 -(-6) det (AT